The coercive force for a certain permanent magnet is 4.0 × 10^4 A/m. This magnet is placed inside a long solenoid of 40 turns/cm and a current is passed in the solenoid to demagnetize it completely. Find the current.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. B = μ base 0 ni, H = B/μ base 0 Given n = 40 turn/cm = 4000 turns/m ⇒ H = ni H = 4 × 10^4 A/m ⇒ i = H/n = 4x 10^10 / 4000 = 10 A.